OpenStudy (anonymous):

Suppose f (x) is a differentiable function defined on the interval [0, 60]. Use the function values given in the table below to answer the following:

OpenStudy (anonymous):

x 0 15 30 45 60 f(x) -3 -1 2 6 2

OpenStudy (anonymous):

a) Approximate the value of f '(20). b) Show that f '(c1) = 1/12 for some value x = c1 in the interval (0, 60). c) Show that f (c2) = 0 for some value x = c2 in the interval (0, 60). Justify your answers.

OpenStudy (anonymous):

@xapproachesinfinity @zepdrix @Hero

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

not sure exactly, but for \(f'(20)\) you could take the slope between the two nearest points \((15,-1),(30,2)\)

OpenStudy (anonymous):

That equals 5 which doesn't seem right?

OpenStudy (anonymous):

it certainly isn't five

OpenStudy (anonymous):

the slope between those two points i mean, definitely not 5

OpenStudy (anonymous):

maybe \(\frac{1}{5}\)?

OpenStudy (anonymous):

Ohhhh oops I flipped it :)

OpenStudy (anonymous):

Do you have any idea how I would figure out b. and c.?? :)

OpenStudy (anonymous):

my guess is a straight forward application of the mean value theorem

OpenStudy (anonymous):

what is the slope between the two points \((0,-3)\) and \((60,2)\) ?

OpenStudy (anonymous):

5/60=1/12?

OpenStudy (anonymous):

lol no, not \(\frac{1}{2}\) ? but rather \[\huge \frac{1}{12}!\]

OpenStudy (anonymous):

and that does it for you if you are familiar with the mean value theorem did you get to that yet?

OpenStudy (anonymous):

Yepp :)

OpenStudy (anonymous):

What would I do for c.?? :)

OpenStudy (anonymous):

function is continuous means it cannot skip values it is negative somewhere, positive somewheres else, that means it has to take on every value between the positive and negative ones that includes 0

OpenStudy (anonymous):

How would I find where it = 0? :)

OpenStudy (anonymous):

somewhere between where it is \(-1\) and \(2\)

OpenStudy (anonymous):

more than this you cannot know

OpenStudy (anonymous):

So could I just say somewhere between f(x)=-1 and 2?

OpenStudy (anonymous):

"somewhere" means the x values somewhere between 15 and 30

OpenStudy (anonymous):

Oh okay, do you think that would be it for the answer then? :)

OpenStudy (anonymous):

Thanks for all your help @satellite73 ! :)